find the product
(y – 10)^2

Answer by lenny460(1073)

(y - 10)^2
(y - 10)(y - 10)
Multiply using the FOIL Method:
y * y = y^2
y * -10 = -10y
-10 * y = -10y
-10 * -10 = 100
y^2 - 10y - 10y + 100
Combine like terms:
-10y - 10y = -20y
The Answer:
y^2 - 20y + 100